# Superposition Prompting: Improving and Accelerating Retrieval-Augmented Generation

### v1

@@ Abstract
Despite the successes of large language models (LLMs), they exhibit significant drawbacks, particularly when processing long contexts. Their inference cost scales quadratically with respect to sequence length, making it expensive for deployment in some real-world text processing applications, such as retrieval-augmented generation (RAG). Additionally, LLMs also exhibit the “distraction phenomenon,” where irrelevant context in the prompt degrades output quality. To address these drawbacks, we propose a novel RAG prompting methodology, superposition prompting, which can be directly applied to pre-trained transformer-based LLMs without the need for fine-tuning. At a high level, superposition prompting allows the LLM to process input documents in parallel prompt paths, discarding paths once they are deemed irrelevant. We demonstrate the capability of our method to simultaneously enhance time efficiency across a variety of question-answering benchmarks using multiple pre-trained LLMs. Furthermore, our technique significantly improves accuracy when the retrieved context is large relative the context the model was trained on. For example, our approach facilitates an 93×93$\times$93 × reduction in compute time while improving accuracy by $43\%$ on the NaturalQuestions-Open dataset with the MPT-7B instruction-tuned model over naive RAG.
@@ 1 Introduction
Figure 2: Comparison of superposition prompting vs. the “classical” (Naive LLM-RAG) prompting paradigm. Squares represents a token, and arrows depict attention dependencies. Whereas the classical approach is a “linked-list” style DAG, superposition prompting arranges token dependencies such that all documents are processed independently. Due to this dependency structure, we can easily leverage the LLM logits to prune irrelevant context, improving long context reasoning. The dependency structure also allows for faster prompt processing, due to the new opportunities for caching and parallelism of the KV cache and logit computations (each gray box represents a logical “batch” that is processed by the LLM, reusing upstream KV caches).
@@
Retrieval-augmented generation (RAG) is one alluring application of transformer-based LLMs. In this setting, the LLM can ground its responses in auxiliary, relevant context. Often, the retrieved documents contain long-form text, leading to the aforementioned downsides (Gao et al., 2023). To improve and accelerate RAG, we propose superposition prompting 111 We drew inspiration from the “path integral” formulation of quantum mechanics (Feynman, 1965), where a particle’s dynamics can be represented as a weighted sum over possible trajectories. Analogously, the language dynamics of superposition prompting are modeled by a weighted sum of possible “token trajectories.” . Superposition prompting is demonstrated to simultaneously improve model accuracy and compute time efficiency for RAG-based question answering tasks without any additional training or finetuning. To highlight our results, we refer the reader to Figure 1.
@@
We plan to release an open-source code implementation.
@@ 2 Related Work
Efficient Long Context Processing. There has been enormous work to reduce the memory footprint and computational costs of transformers using techniques such as compression and KV-caching (Sheng et al., 2023; Lin et al., 2023; Xiao et al., 2022). More efficient versions of the transformer architecture have also been explored. For example, Longformer (Beltagy et al., 2020) introduced a drop-in replacement to the standard self-attention, which makes it scale linearly with sequence length. Similarly, Reformer (Kitaev et al., 2020) uses locality-sensitive hashing to reduce the complexity of attention and improve its efficiency for long sequences. In parallel, the SparseTransformer (Child et al., 2019) focuses on the sparsity of the attention layers. While the above innovation addresses the efficiency of long context processing, they often require non-trivial architecture changes and/or re-training. This makes them impractical to use with existing, pre-trained LLMs (Touvron et al., 2023a; Zhang et al., 2022). Closer to our work are efficient methods which optimize KV caching and consider token importance (Zhang et al., 2023; Liu et al., 2023c). Other works (orthogonal to ours) investigate how to improve efficiency of LLM output generation (Ning et al., 2023). The above methods differ from ours as they investigate acceleration for LLMs generally, whereas we aim to leverage the specifics of the RAG setting to achieve further improvements.
@@ 3 Proposed Method
The retrieval-augmented generation task is comprised of distinct text segments—the preamble (a.k.a. system prompt), a static corpus of documents, and a dynamic, user-supplied query. Instead of concatenating these text segments in textual space, we group them into separate “batches” (the gray boxes in Figure 2), which are passed as calls to the LLM (re-using the KV caches from upstream token segments). With a query as input, superposition prompting processes all choices of documents paired with the query independently (conditioned on the preamble)—in Figure 2, this can be seen as the branching structure. Once the query batch is processed, we then employ path pruning (Section 3.2.3) to discard entire attention dependencies based on an importance metric (the scissors in Figure 2). Both of these optimizations improve inference efficiency and enable the model to discard distracting documents unrelated to the query.
Powered by the structure imposed by our superposition prompting approach, we then propose techniques to further accelerate the inference. First, the high-level notion of token sharing across prompts allows us to employ prompt path caching (Section 3.3.1). Finally, we describe a prompt path parallelization (Section 3.3.2) strategy that leverages independence across segments.
@@ 3.1 Retrieval Augmented Generation
We stylize token sequences as bolded vectors and use $\oplus$ to denote concatenation along the sequence dimension. Supposing there are $n_{d}$ (pre-tokenized) offline documents available for retrieval, we define the set of document token sequences $\{\boldsymbol{d}_{1},\ldots,\boldsymbol{d}_{n_{d}}\}$. We denote the user query as $\boldsymbol{q}$, and our custom preamble sequence as $\boldsymbol{p}$. The goal is to return some response $\boldsymbol{r}$ which answers the query, all while minimizing the latency between the arrival of the query and the serving of the response as observed by the client. The obvious baseline solution (which we refer to as “Naive LLM-RAG”) is where one simply concatenates the input sequences as $\boldsymbol{x}=\boldsymbol{p}\oplus\boldsymbol{d}_{1}\oplus\ldots\oplus\boldsymbol{d}_{n_{d}}\oplus\boldsymbol{q}$, then autoregressively generates $\boldsymbol{r}$ using $\boldsymbol{x}$ as the prompt. However, as shown in Section 4, our approach massively outperforms such a baseline both in terms of quality and performance.
@@ 3.2 Superposition Prompting
We now detail superposition prompting, a new paradigm for prompting language models. In superposition prompting, prompts are not represented as a simple sequence of tokens as they are with “classical” prompting methods (e.g. Naive LLM-RAG). Rather, superpositioned prompts are directed acyclic graphs (DAGs) where nodes are token sequences, and edges codify attention dependencies. Plainly put, a particular token sequence $\boldsymbol{v}$ attends to the tokens in $\boldsymbol{u}$ if there is a path from $\boldsymbol{u}$ to $\boldsymbol{v}$ in the DAG. In this sense, superposition prompting is a generalization of “classical” prompting (since a “classical” prompt is the linked-list special case). Please refer to Algorithm 3 for an algorithmic formalization.
@@ 3.2.1 The ForkJoin Prompt Path Topology
In order to leverage superposition prompting for RAG, we must construct a graph topology out of our text segments. We propose the “ForkJoin” graph structure, depicted in Figure 2. Note that each $\boldsymbol{q}_{i}$ sequence is a duplicate of the original $\boldsymbol{q}$ (Section 3.2.3 will justify this decision). Although this duplication ends up increasing the number of tokens processed, our ablation analysis in Section B.1 demonstrates the superiority of this approach in terms of accuracy. Furthermore, Section 3.3.2 describes how the cost of this duplication can be entirely hidden from the user. The ForkJoin topology (implicitly) results in a pseudo-“local attention” structure (Figure 3). We emphasize that this resulting attention pattern is a construct for visualization only — in reality, all calls to the LLM use fully dense attention, although on relatively smaller context lengths. Concretely, each of the dashed boxes in Figure 2 is a separate LLM call.
@@ 3.2.2 Token Position Assignment
One simple approach could be to truncate token sequences to a common length to enforce a shared positioning. However, truncating may result in loss of input signal if the trimmed tokens contained valuable information for the query. Another approach could be to left-align (or right-pad) sequences to a common length (Gim et al., 2023). While this left aligned padding approach is simple, it yields discontinuities in the prompt sequence position assignments (see Appendix D for quantification). With the ALiBi encoding scheme, these discontinuities will unfairly assign attention penalties to the tokens in shorter documents, since those tokens will be concentrated at earlier token positions 222 An analogous bias would exist if using a right aligned strategy, except shorter documents would be unfairly assigned attention boosts over longer documents. . Thus, we are motivated to propose a positional assignment strategy that does not result in discontinuities.
@@
Note that the path equilibrium positioning strategy results in real-valued positions. This is a departure from common usage of token position assignments, where integer-valued positions are predominant. In Section B.4 we outline why it is reasonable to assume LLMs posses this real-valued positioning capability. We note that the choice of position assignment scheme has no effect on inference efficiency, but can impact model output quality. In Section 5.1, we validate the superiority of path equilibrium positioning.
@@ 3.2.3 Path Pruning
In order to prune paths, we must compute a saliency score for each path. Inspired by SGPT (Muennighoff, 2022), we apply Bayes rule to the output of the language modeling head to compute a saliency score. At a high level, we leverage Bayes’ theorem to compute the posterior distribution
@@
as a saliency metric of document $\boldsymbol{d}_{i}$’s relevancy to the query.333 This resembles the “principle of least action,” which determines the optimal path weighting in the path integral formulation of quantum mechanics. In our experiments, we decide which path indices to prune by greedily sampling the top-$k$ of this categorical distribution (we perform ablations with respect to the choice of $k$ in Section 4.1.2 and Section A.2).
To “physically” apply the pruning, we simply discard the KV caches corresponding to the documents and queries along those paths. Conveniently, all remaining KV caches can be simply concatenated together for use in autoregressive generation of the response.
@@ 3.3.1 Path Caching
Concretely, the preamble KV cache and document KVs are not conditioned on the query, and thus can be precomputed during a “preprocessing” stage. Then, during the “online serving” stage, we retrieve the preamble and document KVs from storage instead of the original input token sequences. Figure 2(d) shows how, during the online serving stage, much of the attention dependencies have already been computed. Note that the memory requirement for employing path caching is a scalar multiple, $c_{\textrm{model}}$, of the raw tokenized sequence length. Here, $c_{\textrm{model}}$ is a fixed scalar that depends on the various aspects of the models, such as number of layers, and embedding dimension (e.g. $c_{\texttt{bloom-7b1}}=492\mathrm{\ KB}$).
@@ 3.3.2 Path Parallelization
Since the superpositioned paths of ForkJoin are independent of each other (by construction), the corresponding KV caches and logits of the query segments can be computed in parallel. While this does not reduce the “CPU time,” it importantly reduces the wall-clock time experienced by the user. The parallelization across the duplicated queries can be accomplished either by (1) concatenating sequences along the batch dimension before inference444 In general, prompt path lengths will vary, thus requiring padding to a common length. However, a length binning strategy may alleviate most overhead in practice. (2) delegating model calls across a distributed cluster of compute nodes (e.g. GPUs), or (3) a combination of batching and distributed inference. The most effective strategy will depend on the specifics of the cluster configuration (e.g. relative network bandwidth vs. available compute per node).
@@ 4 Experimental Results
We perform experiments on two families of large language models, namely BLOOMZ (Muennighoff et al., 2023) and MPT (MosaicML NLP Team, 2023). To quantify the effectiveness of superposition prompting when paired with models of different scales, we use various model sizes from these families. For BLOOMZ, we instantiate 3B parameter (bloomz-3b) and 7.1B parameter (bloomz-7b1) models. For MPT, we use the available instruct fine-tuned 7B parameter model (mpt-7b-instruct). This set of models covers different architectures, sizes, and pretraining recipes. We remind the reader we use the publicly released pretrained checkpoints, without employing any additional training, fine-tuning, or task adaptation.
For our experiments, we are primarily interested in the compute time vs. accuracy tradeoff.555 In our timing and speedup analysis, we follow previous works (Gim et al., 2023) and do not consider the data retrieval portion of the RAG pipeline, since there would be too many assumptions required. We use the fvcore (facebookresearch, 2024) package to compute theoretical floating point operation (FLOP) counts for various inference settings. We evaluate the compute cost of each method in units of compute cycles—similar to FLOPs, but accounting for parallelism. In practice, to achieve the speedups, extra resources (auxiliary memory and/or auxiliary compute for parallelization) will be required. However, as stated, the goal of this exploration is acceleration, not necessarily FLOPs reduction. We refer the reader to Section A.3 for detailed breakdowns of the theoretical speedup gains enabled by each of our proposed optimizations.
@@ 4.1 Results
We leverage the publicly available NaturalQuestions-Open (Liu et al., 2023a) and MuSiQue (Trivedi et al., 2022) datasets. We do not perform any manual prompt tuning or prompt engineering for any method or baseline, and use the same prompts across all experiments (per dataset) to control for discrepancies that could arise with varying prompt wording. For reproducibility, we present the exact prompt wording used for each dataset in Appendix E. We use greedy autoregressive decoding in all experiments, and randomize the order of documents to prevent any systematic bias possible due to location of the “gold documents” (Liu et al., 2023a).

### v2

@@ Abstract
Despite the successes of large language models (LLMs), they exhibit significant drawbacks, particularly when processing long contexts. Their inference cost scales quadratically with respect to sequence length, making it expensive for deployment in some real-world text processing applications, such as retrieval-augmented generation (RAG). Additionally, LLMs also exhibit the “distraction phenomenon”, where irrelevant context in the prompt degrades output quality. To address these drawbacks, we propose a novel RAG prompting methodology, superposition prompting, which can be directly applied to pre-trained transformer-based LLMs without the need for fine-tuning. At a high level, superposition prompting allows the LLM to process input documents in parallel prompt paths, discarding paths once they are deemed irrelevant. We demonstrate the capability of our method to simultaneously enhance time efficiency across a variety of question-answering benchmarks using multiple pre-trained LLMs. Furthermore, our technique significantly improves accuracy when the retrieved context is large relative the context the model was trained on. For example, our approach facilitates a 93×93$\times$93 × reduction in compute time while improving accuracy by $43\%$ on the NaturalQuestions-Open dataset with the MPT-7B instruction-tuned model over naive RAG.
@@ 1 Introduction
Figure 2: Comparison of superposition prompting vs. the “classical” (Naive LLM-RAG) prompting paradigm. Squares represents a token, and arrows depict attention dependencies. Whereas the classical approach is a “linked-list” style DAG, superposition prompting arranges token dependencies such that all documents are processed independently. Due to this dependency structure, we can easily leverage the LLM logits to prune irrelevant context, improving long context reasoning. The dependency structure also allows for faster prompt processing, due to the new opportunities for caching and parallelism of the KV cache and logit computations (each gray box represents, logically, a “batch” that is processed by the LLM, reusing upstream KV caches).
@@
Retrieval-augmented generation (RAG) is one alluring application of transformer-based LLMs. In this setting, the LLM can ground its responses in auxiliary, relevant context. Often, the retrieved documents contain long-form text, leading to the aforementioned downsides (Gao et al., 2023). To improve and accelerate RAG, we propose superposition prompting 111 We drew inspiration from the “path integral” formulation of quantum mechanics (Feynman, 1965), where a particle’s dynamics can be represented as a weighted sum over possible trajectories. Analogously, the language dynamics of superposition prompting are modeled by a weighted sum of possible “token trajectories”. . Superposition prompting is demonstrated to simultaneously improve model accuracy and compute time efficiency for RAG-based question answering tasks without any additional training or fine-tuning. To highlight our results, we refer the reader to Figure 1.
@@
For reproducibility, our implementation can be found at https://github.com/apple/ml-superposition-prompting.
@@ 2 Related Work
Efficient Long Context Processing. There have been significant efforts to reduce the memory footprint and computational costs of transformers using techniques such as compression and KV-caching (Sheng et al., 2023; Lin et al., 2023; Xiao et al., 2022). More efficient versions of the transformer architecture have also been explored. For example, Longformer (Beltagy et al., 2020) introduced a drop-in replacement to the standard self-attention, which makes it scale linearly with sequence length. Similarly, Reformer (Kitaev et al., 2020) uses locality-sensitive hashing to reduce the complexity of attention and improve its efficiency for long sequences. In parallel, the SparseTransformer (Child et al., 2019) focuses on the sparsity of the attention layers. While the above innovation addresses the efficiency of long context processing, they often require non-trivial architecture changes and/or re-training. This makes them impractical to use with existing, pre-trained LLMs (Touvron et al., 2023; Zhang et al., 2022). Closer to our work are efficient methods which optimize KV caching and consider token importance (Zhang et al., 2023; Liu et al., 2023c). Other works (orthogonal to ours) investigate how to improve efficiency of LLM output generation (Ning et al., 2023). The above methods differ from ours as they investigate acceleration for LLMs generally, whereas we aim to leverage the specifics of the RAG setting to achieve further improvements.
@@ 3 Proposed Method
The retrieval-augmented generation task is comprised of distinct text segments—the preamble (a.k.a. system prompt), a (static) corpus of documents, and a (dynamic) user-supplied query. Instead of concatenating these text segments in textual space, we group them into separate “batches” (the gray boxes in Figure 2), which are passed as calls to the LLM (re-using the KV caches from upstream token segments). With a query as input, superposition prompting processes all choices of documents paired with the query independently (conditioned on the preamble)—in Figure 2, this can be seen as the branching structure. Once the query batch is processed, we then employ path pruning (Section 3.2.3) to discard entire attention dependencies based on an importance metric (the scissors in Figure 2). Both of these optimizations improve inference efficiency and enable the model to discard distracting documents unrelated to the query.
Enabled by the added structure of our superposition prompting approach, we then propose techniques to further accelerate the inference. First, the high-level notion of token sharing across prompts allows us to employ prompt path caching (Section 3.3.1). Finally, we describe a prompt path parallelization (Section 3.3.2) strategy that leverages independence across segments.
@@ 3.1 Retrieval Augmented Generation
We stylize token sequences as bolded vectors and use $\oplus$ to denote concatenation along the sequence dimension. Supposing there are $n_{d}$ (pre-tokenized) offline documents available for retrieval, we define the set of document token sequences $\{\boldsymbol{d}_{1},\ldots,\boldsymbol{d}_{n_{d}}\}$. We denote the user query as $\boldsymbol{q}$, and our custom preamble sequence as $\boldsymbol{p}$. The goal is to return some response $\boldsymbol{r}$ which answers the query, all while minimizing the latency between the arrival of the query and the serving of the response as observed by the client. The obvious baseline solution (which we refer to as “Naive LLM-RAG”) is where one simply concatenates the input sequences as $\boldsymbol{x}=\boldsymbol{p}\oplus\boldsymbol{d}_{1}\oplus\cdots\oplus\boldsymbol{d}_{n_{d}}\oplus\boldsymbol{q}$, then autoregressively generates $\boldsymbol{r}$ using $\boldsymbol{x}$ as the prompt. However, as shown in Section 4, our approach massively outperforms such a baseline both in terms of quality and performance.
@@ 3.2 Superposition Prompting
We now detail superposition prompting, a new paradigm for prompting language models. In superposition prompting, prompts are not represented as a simple sequence of tokens as they are with “classical” prompting methods (e.g. Naive LLM-RAG). Rather, superpositioned prompts are directed acyclic graphs (DAGs) where nodes are token sequences, and edges codify attention dependencies. Plainly put, a particular token sequence, $\boldsymbol{v}$, attends to the tokens in another token sequence, $\boldsymbol{u}$, if and only if there is a path from $\boldsymbol{u}$ to $\boldsymbol{v}$ in the DAG. In this sense, superposition prompting is a generalization of “classical” prompting (since a “classical” prompt is the linked-list special case). Please refer to Algorithm 3 for an algorithmic formalization.
@@ 3.2.1 The ForkJoin Prompt Path Topology
In order to leverage superposition prompting for RAG, we must construct a graph topology out of our text segments. We propose the ForkJoin graph structure, depicted in Figure 2. Note that each $\boldsymbol{q}_{i}$ sequence is a duplicate of the original $\boldsymbol{q}$ (Section 3.2.3 will justify this decision). Although this duplication ends up increasing the number of tokens processed, our ablation analysis in Section B.1 demonstrates the superiority of this approach in terms of accuracy. Furthermore, Section 3.3.2 describes how the cost of this duplication can be entirely hidden from the user. The ForkJoin topology (implicitly) results in a pseudo-“local attention” structure (Figure 3). We emphasize that this resulting attention pattern is a construct for visualization only—in reality, all calls to the LLM use fully dense attention, although on relatively smaller context lengths. Concretely, each of the dashed boxes in Figure 2 is a separate LLM call.
@@ 3.2.2 Token Position Assignment
One simple approach could be to truncate token sequences to a common length to enforce a shared positioning. However, truncating may result in loss of input signal if the trimmed tokens contained valuable information for the query. Another approach could be to left-align (or right-pad) sequences to a common length (Gim et al., 2023). While this left aligned padding approach is simple, it yields discontinuities in the prompt sequence position assignments (see Appendix E for quantification). With the ALiBi encoding scheme (Press et al., 2021), it can be easily shown that discontinuities unfairly assign attention penalties to the tokens in shorter documents, since the tokens will be concentrated at earlier token positions (and thus larger distances from the current token).222 An analogous bias would exist if using a right aligned strategy, except shorter documents would be unfairly assigned attention boosts over longer documents. Thus, we are motivated to propose a positional assignment strategy that does not result in discontinuities.
@@
Note that the path equilibrium positioning strategy results in real-valued positions. This is a departure from common usage of token position assignments, where integer-valued positions are predominant.333 While this is trivial for the ALiBi positional encoding, it is non-trivial for the Rotary Position Embedding (Su et al., 2021) scheme. To handle this case, we linearly interpolate the argument of the sinusoidal functions. We note that the choice of position assignment scheme has no effect on inference efficiency, but can impact model output quality. In Section 5.1, we validate the superiority of path equilibrium positioning.
@@ 3.2.3 Path Pruning
In order to prune paths, we must compute a saliency score for each path. Inspired by SGPT (Muennighoff, 2022), we apply Bayes rule to the output of the language modeling head to compute a saliency or entailment score. At a high level, we leverage Bayes’ theorem to compute the posterior distribution
@@
as a saliency metric of document $\boldsymbol{d}_{i}$’s relevancy to the query.444 This resembles the “principle of least action,” which determines the optimal path weighting in the path integral formulation of quantum mechanics. In our experiments, we decide which path indices to prune by greedily selecting the top-$k$ of this categorical distribution (we perform ablations with respect to the choice of $k$ in Section 4.1.2 and Section A.1).
To “physically” apply the pruning, we can simply discard the KV caches corresponding to the documents and queries along those paths. Conveniently, all remaining KV caches can be simply concatenated together for use in autoregressive generation of the response.
@@ 3.3.1 Path Caching
We now describe our path caching mechanism. Concretely, the preamble KV cache and document KVs are not conditioned on the query, and thus can be precomputed during a “preprocessing” stage. Then, during the “online serving” stage, we retrieve the preamble and document KVs from storage instead of the original input token sequences. Figure 3(d) shows how, during the online serving stage, much of the attention dependencies have already been computed. Note that the memory requirement for employing path caching is a scalar multiple, $c_{\textrm{model}}$, of the raw tokenized sequence length. Here, $c_{\textrm{model}}$ is a fixed scalar that depends on the various aspects of the models, such as number of layers, and embedding dimension (e.g. $c_{\texttt{bloom-7b1}}=492\mathrm{\ KB}$).
@@ 3.3.2 Path Parallelization
Since the superpositioned paths of ForkJoin are independent of each other (by construction), the corresponding KV caches and logits of the query segments can be computed in parallel. While this does not reduce the “CPU time,” it importantly reduces the wall-clock time experienced by the user. The parallelization across the duplicated queries can be accomplished either by (1) concatenating sequences along the batch dimension before inference555 In general, prompt path lengths will vary, thus requiring padding to a common length. However, a length binning strategy may alleviate most overhead in practice. (2) delegating model calls across a distributed cluster of compute nodes (e.g. GPUs), or (3) a combination of batching and distributed inference. The most effective strategy will depend on the specifics of the cluster configuration (e.g. relative network bandwidth vs. available compute per node).
@@ 4 Experimental Results
We perform experiments on three families of large language models, namely OpenELM (Mehta et al., 2024), BLOOMZ (Muennighoff et al., 2023), and MPT (MosaicML NLP Team, 2023). To quantify the effectiveness of superposition prompting when paired with models of different scales, we use various model sizes from these families. For OpenELM, we use the 3B-Instruct configuration. For BLOOMZ, we instantiate 3B parameter (bloomz-3b) and 7.1B parameter (bloomz-7b1) models. Finally, for MPT, we use the available instruct fine-tuned 7B parameter model (mpt-7b-instruct). This set of models covers different architectures, positional encoding schemes, sizes, and pretraining recipes. We remind the reader we use the publicly released pretrained checkpoints, without employing any additional training, fine-tuning, or task adaptation.
For our experiments, we are primarily interested in the compute time vs. accuracy tradeoff.666 In our timing and speedup analysis, we follow previous works (Gim et al., 2023) and do not consider the data retrieval portion of the RAG pipeline, which would require too many assumptions. We use the fvcore (facebookresearch, 2024) package to compute theoretical floating point operation (FLOP) counts for various inference settings. We evaluate the compute cost of each method in units of compute cycles—similar to FLOPs, but accounting for parallelism. In practice, to achieve the speedups, extra resources (auxiliary memory and/or auxiliary compute for parallelization) will be required. However, as stated, the goal of this exploration is acceleration, not necessarily FLOPs reduction. We refer the reader to Section A.2 for detailed breakdowns of the theoretical speedup gains enabled by each of our proposed optimizations.
@@ 4.1 Results
We leverage the publicly available NaturalQuestions-Open (Liu et al., 2023a) and MuSiQue (Trivedi et al., 2022) datasets. We do not perform any manual prompt tuning or prompt engineering for any method or baseline, and use the same prompts across all experiments (per dataset) to control for discrepancies that could arise with varying prompt wording. For reproducibility, we present the exact prompt wording used for each dataset in Appendix F. We use greedy autoregressive decoding in all experiments, and randomize the order of documents to prevent any systematic bias possible due to location of the “gold documents” (à la Liu et al., 2023a).